OpenStudy (anonymous):

Solve for y. xy+2x=8

OpenStudy (anonymous):

x(y+2)=8 y+2=8/x y=(8/x)-2

OpenStudy (anonymous):

xy + 2x = 8 (divide everything by x) y + 2 = 8/x -2 -2 y = 8/x - 2